# A Survey of Reinforcement Learning For Economics

### 17 Mar 2026 v3

@@ 9.1.1 The Stochastic Game Framework
An $n$-player stochastic game $\Gamma=(n,\mathcal{S},\mathcal{A}_{1},\ldots,\mathcal{A}_{n},P,r_{1},\ldots,r_{n},\gamma)$ consists of a finite state space $\mathcal{S}$; finite action sets $\mathcal{A}_{i}$ for each player $i$; a transition function $P:\mathcal{S}\times\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{n}\to\Delta(\mathcal{S})$; reward functions $r_{i}:\mathcal{S}\times\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{n}\to\mathbb{R}$; and a common discount factor $\gamma\in[0,1)$. At each stage, all players simultaneously choose actions, receive individual rewards, and the game transitions to a new state.159159159Shapley (1953) introduced stochastic games in 1953 for the two-player zero-sum case, proving existence of the value via a contraction argument on the Bellman operator. The general-sum extension to $n$ players is due to Fink (1964) and Takahashi (1964).
@@ 9.1.2 Minimax-Q Learning
Minimax-Q converges to the minimax Q-values under Robbins-Monro learning rates ($\sum_{t}\alpha_{t}=\infty$, $\sum_{t}\alpha_{t}^{2}<\infty$) and infinite exploration of all state-action tuples.160160160The convergence proof extends the contraction argument for standard Q-learning. Because the zero-sum minimax operator is a contraction with modulus $\gamma$ under the $\ell^{\infty}$ norm, the stochastic approximation converges to the fixed point. See Littman (1994) and the general treatment in Szepesvári and Littman (1999). However, the algorithm sacrifices rationality: it plays the equilibrium strategy even against exploitable opponents.161161161In the soccer game of Littman (1994), minimax-Q won 53.7% against a hand-built opponent versus 26.1% for Q-learning. Against an adversarial challenger, Q-learning won 0% (its deterministic policy was fully predictable); minimax-Q won 37.5% through mixed strategies.
@@ 9.1.3 Nash-Q Learning
These conditions are restrictive. When stage games have multiple Nash equilibria, agents may select different equilibria for their backups, causing divergence.162162162In the experiments of Hu and Wellman (2003), a grid game with a unique equilibrium Q-function converged in 100% of trials under Nash-Q versus 20% under independent Q-learning; a game with three equilibrium Q-functions converged in only 68–90% of trials. Nash-Q requires each agent to observe all other agents’ rewards and to maintain Q-values over the joint action space $\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{n}$, with storage $O(n|\mathcal{S}|\prod_{i}|\mathcal{A}_{i}|)$, exponential in the number of players. Nash-Q reduces to standard Q-learning in the single-agent case.
@@ 9.1.4 The Convergence Problem
Two further algorithms deserve mention. Friend-or-Foe Q-learning (Littman, 2001) decomposes agents as cooperative (friend) or adversarial (foe), using $\max$ for friends and minimax for foes; it always converges but requires knowing the relationship type a priori.163163163Correlated-Q learning (Greenwald and Hall, 2003) generalizes both Nash-Q and Minimax-Q by using correlated equilibrium, a probability distribution over joint actions enforced by a correlating device. The set of correlated equilibria contains all Nash equilibria. Correlated-Q converges under conditions analogous to Nash-Q. The evolutionary perspective of Börgers and Sarin (1997) provides a deeper lens: reinforcement learning dynamics in matrix games converge to the replicator equation from evolutionary game theory. The cycling of Q-learning in games such as matching pennies is structurally identical to the cycling of replicator dynamics in Rock-Paper-Scissors games.
@@ 9.1.5 Simulation Study: Cournot and Bertrand Duopoly
Two canonical games from industrial organization serve as benchmarks. In Cournot duopoly, two firms choose quantities $q_{i}\in\{0,1,\ldots,9\}$ with inverse demand P(Q)=10QP(Q)=10-Q and marginal cost $c=1$; the unique Nash equilibrium is $q^{*}=3$ with profit $\pi^{*}=9$. In Bertrand duopoly with differentiated products, two firms choose prices $p_{i}\in\{0,1,\ldots,9\}$ with demand $d_{i}=10-2p_{i}+p_{j}$ and marginal cost $c=1$; the continuous Nash equilibrium is $p^{*}\approx 4.33$, which discretizes to $p^{*}=4$ with profit $\pi^{*}=18$. Both games have unique Nash equilibria in pure strategies.164164164Each configuration runs for 50,000 iterations across 20 seeds. Three algorithms compete: independent Q-learning (IQL), Nash-Q, and WoLF-PHC.
@@ 9.2 Counterfactual Regret Minimization
where $\pi^{\sigma}_{-i}(h)$ is the probability opponents reach $h$, and $\pi^{\sigma}(ha,z)$ is the probability of reaching terminal $z$ from $ha$. The counterfactual formulation weights by opponent reach $\pi^{\sigma}_{-i}(h)$ rather than joint reach $\pi^{\sigma}(h)$, ensuring non-zero updates even at rarely-visited information sets. Cumulative regret for action $a$ is $R^{T}(I,a)=\sum_{t=1}^{T}[v^{\sigma^{t}}(I,a)-v^{\sigma^{t}}(I)]$. CFR updates via regret matching:165165165Regret matching is an action selection rule where the probability of choosing action $a$ is proportional to the cumulative regret for not having chosen $a$ in the past (truncated at zero). Actions with high regret receive higher probability; actions with negative regret (having performed worse than average) receive zero probability.
@@
The average strategy $\bar{\sigma}^{T}$ converges to $\varepsilon$-Nash with $\varepsilon=O(|\mathcal{I}|\sqrt{|\mathcal{A}|}\Delta/\sqrt{T})$, where $\Delta$ is the range of payoffs (Zinkevich et al., 2008).166166166An $\varepsilon$-Nash equilibrium is a strategy profile where no player can improve her expected payoff by more than $\varepsilon$ through unilateral deviation. As $\varepsilon\to 0$, this converges to an exact Nash equilibrium.
CFR+ (Tammelin, 2014) replaces standard regret matching with Regret Matching+, which truncates cumulative regrets to zero after each update rather than only at action selection. The update becomes $R^{T+1}(I,a)=\max(R^{T}(I,a)+r^{T+1}(I,a),\,0)$, where $r^{T+1}(I,a)=v^{\sigma^{T+1}}(I,a)-v^{\sigma^{T+1}}(I)$ is the instantaneous counterfactual regret. CFR+ also weights iteration $t$ by $t$ when computing the average strategy. While vanilla CFR converges at $O(1/\sqrt{T})$, CFR+ empirically converges at O(1/T)O(1/T).167167167The $O(1/T)$ rate for CFR+ is empirically observed but not yet proven in full generality. Tammelin (2014) conjectured this rate based on extensive experiments across poker variants. CFR+ enabled Bowling et al. (2015) to essentially solve heads-up limit Texas hold’em, a game with $3.16\times 10^{17}$ states, the first non-trivial imperfect-information game played competitively by humans to be essentially solved. Their program Cepheus achieved exploitability below 1 mbb/g (milli-big-blind per game).
@@ 9.3.2 Neural Fictitious Self-Play
NFSP (Heinrich and Silver, 2016) combines fictitious play168168168Fictitious play (Brown, 1951) is a classical learning rule where each player best-responds to the empirical distribution of opponents’ past actions. Under certain conditions (e.g., two-player zero-sum games), the time-average strategies converge to Nash equilibrium. (Brown, 1951) with deep Q-learning.169169169Deep Q-Network (DQN) (Mnih et al., 2015) approximates the Q-function with a neural network, using experience replay and a target network to stabilize training. The target network $\theta^{-}$ in the loss function is a delayed copy of the main network, updated periodically. Each player maintains two networks: a best-response network $Q_{\theta}$ trained via DQN, and an average strategy network $\bar{\pi}_{\phi}$ trained via supervised learning on the best-response actions:
@@ 9.3.3 Poker Results
These methods achieved superhuman performance in poker. Deep CFR attained exploitability of 37 mbb/g170170170Exploitability measures how far a strategy is from Nash equilibrium, defined as the maximum expected gain an adversary could achieve by best-responding. Zero exploitability means the strategy is unexploitable (Nash). In poker, exploitability is measured in milli-big-blinds per game (mbb/g). in heads-up flop hold’em (Brown et al., 2019). Libratus defeated top human professionals in heads-up no-limit hold’em using nested subgame solving with CFR (Brown and Sandholm, 2018). Pluribus extended to six-player no-limit hold’em (Brown and Sandholm, 2019). The game tree of heads-up no-limit hold’em contains $\sim 10^{161}$ states; these results demonstrate that CFR-based methods scale to economically relevant game sizes.
@@ 9.4 The Coase Conjecture
Coase (1972) conjectured that a durable goods monopolist171171171A durable good provides utility over multiple periods (e.g., a car, appliance, or software license) rather than being consumed immediately. Unlike non-durable goods, durable goods create intertemporal competition, as the seller at time $t$ competes with her own future self at $t+1$. loses market power when buyers are patient. Unable to commit to future prices, the seller competes with her future self, eroding rents. As the inter-offer interval shrinks ($\delta\to 1$), price collapses to marginal cost and all surplus is extracted by buyers. Gul et al. (1986) formalized this for the gap case, where the seller’s cost is strictly below all buyer valuations, guaranteeing trade and a unique stationary equilibrium.172172172Stokey (1981) showed that a monopolist facing rational consumers cannot price discriminate intertemporally; Bulow (1982) proved that in the no-gap case, the monopolist prefers renting to selling. The no-gap case admits multiple equilibria with more complex dynamics.
@@ 9.4.2 Equilibrium Analysis
For $\pi<\pi^{*}$, the seller pools (offers $v_{L}$ immediately). For $\pi>\pi^{*}$, the seller screens (offers $P^{*}$, then $v_{L}$ if rejected).173173173In a screening equilibrium, the seller uses price to separate buyer types: high-value buyers accept immediately at a high price, while low-value buyers reject and receive a lower offer. In a pooling equilibrium, the seller offers a single price that all types accept. The gap case ($0=c<v_{L}$) guarantees trade in equilibrium. The Coase conjecture manifests as $\delta\to 1$, where $P^{*}(\delta)\to v_{L}$ and the seller cannot extract surplus from high types.
@@ 9.4.3 Computational Results
I model the bargaining game as an extensive-form game and apply CFR.174174174The game tree has two information sets for the seller (indexed by rejection history) and two for the buyer (indexed by private type). CFR runs for 5,000 iterations per parameter configuration; NashConv (sum of exploitabilities) measures convergence.
@@ 10 Bandits and Dynamic Pricing
This chapter focuses on online reinforcement learning (Section 2), where, unlike the simulator-based training of preceding chapters, the agent learns directly from interactions with real customers and exploration is a real cost that must be balanced against exploitation. A seller faces $T$ customers in sequence, setting a price for each and observing only whether the customer bought.176176176Rothschild (1974) posed pricing under demand uncertainty as a two-armed bandit problem. His insight was that a myopic seller can get stuck at a suboptimal price forever, because exploiting the currently best-looking price generates no information about alternatives. The seller does not observe the customer’s willingness to pay. Regret, the total revenue gap between the seller’s policy and the best fixed price in hindsight, is the central measure: if the optimal fixed price earns $r^{*}$ per customer, a policy with regret $R(T)$ earns $Tr^{*}-R(T)$ in total. The central question is how fast $R(T)$ shrinks as the seller accumulates purchase data, and how structural assumptions about demand affect this rate.
@@ 10.1.1 No Structure on Demand
Kleinberg and Leighton (2003) study the simplest version of the problem. Customers arrive with valuations drawn i.i.d. from an unknown distribution on $[0,1]$, and the seller posts a price from a continuous set. The demand curve $D(p)=\Pr(v\geq p)$ is unknown; the only feedback is whether each customer bought. The seller’s goal is to minimize regret against the single price $p^{*}$ that maximizes $p\cdot D(p)$. Kleinberg and Leighton prove that the minimax regret is $\Theta(\sqrt{T})$.177177177The upper bound, $O(\sqrt{T\log T})$, discretizes $[0,1]$ into $K=\lceil(T/\log T)^{1/4}\rceil$ prices and runs the UCB1 algorithm of Auer et al. (2002a). The lower bound, $\Omega(\sqrt{T})$, constructs a family of demand curves parameterized by the location of the optimal price $p^{*}\in[0.3,0.4]$. The key tension: posting prices far from $p^{*}$ is informative about demand but costly in revenue; posting prices near $p^{*}$ is cheap but uninformative. Resolving this tension costs at least $\Omega(\sqrt{T})$ in cumulative revenue. UCB1 (Auer et al., 2002a) selects the price maximizing $\hat{\mu}_{p_{k}}(t)+\sqrt{2\ln t/N_{p_{k}}(t)}$, where $\hat{\mu}_{p_{k}}(t)$ is the empirical mean profit and $N_{p_{k}}(t)$ the number of trials; the second term is an exploration bonus that shrinks as a price is tried more, implementing the principle of optimism in the face of uncertainty. In concrete terms, after 10,000 customers the seller loses roughly 100 customers’ worth of revenue to the uncertainty in demand. No algorithm can do better without imposing structure on $D$. For adversarial valuations (chosen by a worst-case opponent rather than drawn from a fixed distribution), the minimax regret rises to $\Theta(T^{2/3})$, achieved by the Exp3 algorithm178178178Exp3 (Auer et al., 2002b) maintains a weight $w_{k}(t)$ for each price, selecting $p_{k}$ with probability proportional to $w_{k}(t)$ and updating the chosen price’s weight by $\exp(\eta\hat{r}_{k,t})$ where $\hat{r}_{k,t}$ is the revenue importance-weighted by the selection probability; because no model of demand is assumed, the guarantee holds against an adversary who chooses valuations after observing the algorithm. on a discretized price grid. I focus on the stochastic setting throughout this chapter.

### 23 Mar 2026 v4

@@ 9.1.1 The Stochastic Game Framework
An $n$-player stochastic game $\Gamma=(n,\mathcal{S},\mathcal{A}_{1},\ldots,\mathcal{A}_{n},P,r_{1},\ldots,r_{n},\gamma)$ consists of a finite state space $\mathcal{S}$; finite action sets $\mathcal{A}_{i}$ for each player $i$; a transition function $P:\mathcal{S}\times\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{n}\to\Delta(\mathcal{S})$; reward functions $r_{i}:\mathcal{S}\times\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{n}\to\mathbb{R}$; and a common discount factor $\gamma\in[0,1)$. At each stage, all players simultaneously choose actions, receive individual rewards, and the game transitions to a new state.169169169Shapley (1953) introduced stochastic games in 1953 for the two-player zero-sum case, proving existence of the value via a contraction argument on the Bellman operator. The general-sum extension to $n$ players is due to Fink (1964) and Takahashi (1964).
@@ 9.1.2 Minimax-Q Learning
Minimax-Q converges to the minimax Q-values under Robbins-Monro learning rates ($\sum_{t}\alpha_{t}=\infty$, $\sum_{t}\alpha_{t}^{2}<\infty$) and infinite exploration of all state-action tuples.170170170The convergence proof extends the contraction argument for standard Q-learning. Because the zero-sum minimax operator is a contraction with modulus $\gamma$ under the $\ell^{\infty}$ norm, the stochastic approximation converges to the fixed point. See Littman (1994) and the general treatment in Szepesvári and Littman (1999). However, the algorithm sacrifices rationality: it plays the equilibrium strategy even against exploitable opponents.171171171In the soccer game of Littman (1994), minimax-Q won 53.7% against a hand-built opponent versus 26.1% for Q-learning. Against an adversarial challenger, Q-learning won 0% (its deterministic policy was fully predictable); minimax-Q won 37.5% through mixed strategies.
@@ 9.1.3 Nash-Q Learning
These conditions are restrictive. When stage games have multiple Nash equilibria, agents may select different equilibria for their backups, causing divergence.172172172In the experiments of Hu and Wellman (2003), a grid game with a unique equilibrium Q-function converged in 100% of trials under Nash-Q versus 20% under independent Q-learning; a game with three equilibrium Q-functions converged in only 68–90% of trials. Nash-Q requires each agent to observe all other agents’ rewards and to maintain Q-values over the joint action space $\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{n}$, with storage $O(n|\mathcal{S}|\prod_{i}|\mathcal{A}_{i}|)$, exponential in the number of players. Nash-Q reduces to standard Q-learning in the single-agent case.
@@ 9.1.4 The Convergence Problem
Two further algorithms deserve mention. Friend-or-Foe Q-learning (Littman, 2001) decomposes agents as cooperative (friend) or adversarial (foe), using $\max$ for friends and minimax for foes; it always converges but requires knowing the relationship type a priori.173173173Correlated-Q learning (Greenwald and Hall, 2003) generalizes both Nash-Q and Minimax-Q by using correlated equilibrium, a probability distribution over joint actions enforced by a correlating device. The set of correlated equilibria contains all Nash equilibria. Correlated-Q converges under conditions analogous to Nash-Q. The evolutionary perspective of Börgers and Sarin (1997) provides a deeper lens: reinforcement learning dynamics in matrix games converge to the replicator equation from evolutionary game theory. The cycling of Q-learning in games such as matching pennies is structurally identical to the cycling of replicator dynamics in Rock-Paper-Scissors games.
@@ 9.1.5 Simulation Study: Cournot and Bertrand Duopoly
Two canonical games from industrial organization serve as benchmarks. In Cournot duopoly, two firms choose quantities $q_{i}\in\{0,1,\ldots,9\}$ with inverse demand P(Q)=10QP(Q)=10-Q and marginal cost $c=1$; the unique Nash equilibrium is $q^{*}=3$ with profit $\pi^{*}=9$. In Bertrand duopoly with differentiated products, two firms choose prices $p_{i}\in\{0,1,\ldots,9\}$ with demand $d_{i}=10-2p_{i}+p_{j}$ and marginal cost $c=1$; the continuous Nash equilibrium is $p^{*}\approx 4.33$, which discretizes to $p^{*}=4$ with profit $\pi^{*}=18$. Both games have unique Nash equilibria in pure strategies.174174174Each configuration runs for 50,000 iterations across 20 seeds. Three algorithms compete: independent Q-learning (IQL), Nash-Q, and WoLF-PHC.
@@ 9.2 Counterfactual Regret Minimization
where $\pi^{\sigma}_{-i}(h)$ is the probability opponents reach $h$, and $\pi^{\sigma}(ha,z)$ is the probability of reaching terminal $z$ from $ha$. The counterfactual formulation weights by opponent reach $\pi^{\sigma}_{-i}(h)$ rather than joint reach $\pi^{\sigma}(h)$, ensuring non-zero updates even at rarely-visited information sets. Cumulative regret for action $a$ is $R^{T}(I,a)=\sum_{t=1}^{T}[v^{\sigma^{t}}(I,a)-v^{\sigma^{t}}(I)]$. CFR updates via regret matching:175175175Regret matching is an action selection rule where the probability of choosing action $a$ is proportional to the cumulative regret for not having chosen $a$ in the past (truncated at zero). Actions with high regret receive higher probability; actions with negative regret (having performed worse than average) receive zero probability.
@@
The average strategy $\bar{\sigma}^{T}$ converges to $\varepsilon$-Nash with $\varepsilon=O(|\mathcal{I}|\sqrt{|\mathcal{A}|}\Delta/\sqrt{T})$, where $\Delta$ is the range of payoffs (Zinkevich et al., 2008).176176176An $\varepsilon$-Nash equilibrium is a strategy profile where no player can improve her expected payoff by more than $\varepsilon$ through unilateral deviation. As $\varepsilon\to 0$, this converges to an exact Nash equilibrium.
CFR+ (Tammelin, 2014) replaces standard regret matching with Regret Matching+, which truncates cumulative regrets to zero after each update rather than only at action selection. The update becomes $R^{T+1}(I,a)=\max(R^{T}(I,a)+r^{T+1}(I,a),\,0)$, where $r^{T+1}(I,a)=v^{\sigma^{T+1}}(I,a)-v^{\sigma^{T+1}}(I)$ is the instantaneous counterfactual regret. CFR+ also weights iteration $t$ by $t$ when computing the average strategy. While vanilla CFR converges at $O(1/\sqrt{T})$, CFR+ empirically converges at O(1/T)O(1/T).177177177The $O(1/T)$ rate for CFR+ is empirically observed but not yet proven in full generality. Tammelin (2014) conjectured this rate based on extensive experiments across poker variants. CFR+ enabled Bowling et al. (2015) to essentially solve heads-up limit Texas hold’em, a game with $3.16\times 10^{17}$ states, the first non-trivial imperfect-information game played competitively by humans to be essentially solved. Their program Cepheus achieved exploitability below 1 mbb/g (milli-big-blind per game).
@@ 9.3.2 Neural Fictitious Self-Play
NFSP (Heinrich and Silver, 2016) combines fictitious play178178178Fictitious play (Brown, 1951) is a classical learning rule where each player best-responds to the empirical distribution of opponents’ past actions. Under certain conditions (e.g., two-player zero-sum games), the time-average strategies converge to Nash equilibrium. (Brown, 1951) with deep Q-learning.179179179Deep Q-Network (DQN) (Mnih et al., 2015) approximates the Q-function with a neural network, using experience replay and a target network to stabilize training. The target network $\theta^{-}$ in the loss function is a delayed copy of the main network, updated periodically. Each player maintains two networks: a best-response network $Q_{\theta}$ trained via DQN, and an average strategy network $\bar{\pi}_{\phi}$ trained via supervised learning on the best-response actions:
@@ 9.3.3 Poker Results
These methods achieved superhuman performance in poker. Deep CFR attained exploitability of 37 mbb/g180180180Exploitability measures how far a strategy is from Nash equilibrium, defined as the maximum expected gain an adversary could achieve by best-responding. Zero exploitability means the strategy is unexploitable (Nash). In poker, exploitability is measured in milli-big-blinds per game (mbb/g). in heads-up flop hold’em (Brown et al., 2019). Libratus defeated top human professionals in heads-up no-limit hold’em using nested subgame solving with CFR (Brown and Sandholm, 2018). Pluribus extended to six-player no-limit hold’em (Brown and Sandholm, 2019). The game tree of heads-up no-limit hold’em contains $\sim 10^{161}$ states; these results demonstrate that CFR-based methods scale to economically relevant game sizes.
@@ 9.4 The Coase Conjecture
Coase (1972) conjectured that a durable goods monopolist181181181A durable good provides utility over multiple periods (e.g., a car, appliance, or software license) rather than being consumed immediately. Unlike non-durable goods, durable goods create intertemporal competition, as the seller at time $t$ competes with her own future self at $t+1$. loses market power when buyers are patient. Unable to commit to future prices, the seller competes with her future self, eroding rents. As the inter-offer interval shrinks ($\delta\to 1$), price collapses to marginal cost and all surplus is extracted by buyers. Gul et al. (1986) formalized this for the gap case, where the seller’s cost is strictly below all buyer valuations, guaranteeing trade and a unique stationary equilibrium.182182182Stokey (1981) showed that a monopolist facing rational consumers cannot price discriminate intertemporally; Bulow (1982) proved that in the no-gap case, the monopolist prefers renting to selling. The no-gap case admits multiple equilibria with more complex dynamics.
@@ 9.4.2 Equilibrium Analysis
For $\pi<\pi^{*}$, the seller pools (offers $v_{L}$ immediately). For $\pi>\pi^{*}$, the seller screens (offers $P^{*}$, then $v_{L}$ if rejected).183183183In a screening equilibrium, the seller uses price to separate buyer types: high-value buyers accept immediately at a high price, while low-value buyers reject and receive a lower offer. In a pooling equilibrium, the seller offers a single price that all types accept. The gap case ($0=c<v_{L}$) guarantees trade in equilibrium. The Coase conjecture manifests as $\delta\to 1$, where $P^{*}(\delta)\to v_{L}$ and the seller cannot extract surplus from high types.
@@ 9.4.3 Computational Results
I model the bargaining game as an extensive-form game and apply CFR.184184184The game tree has two information sets for the seller (indexed by rejection history) and two for the buyer (indexed by private type). CFR runs for 5,000 iterations per parameter configuration; NashConv (sum of exploitabilities) measures convergence.
@@ 10 Bandits and Dynamic Pricing
This chapter focuses on online reinforcement learning (Section 2), where, unlike the simulator-based training of preceding chapters, the agent learns directly from interactions with real customers and exploration is a real cost that must be balanced against exploitation. A seller faces $T$ customers in sequence, setting a price for each and observing only whether the customer bought.186186186Rothschild (1974) posed pricing under demand uncertainty as a two-armed bandit problem. His insight was that a myopic seller can get stuck at a suboptimal price forever, because exploiting the currently best-looking price generates no information about alternatives. The seller does not observe the customer’s willingness to pay. Regret, the total revenue gap between the seller’s policy and the best fixed price in hindsight, is the central measure: if the optimal fixed price earns $r^{*}$ per customer, a policy with regret $R(T)$ earns $Tr^{*}-R(T)$ in total. The central question is how fast $R(T)$ shrinks as the seller accumulates purchase data, and how structural assumptions about demand affect this rate.
@@ 10.1.1 No Structure on Demand
Kleinberg and Leighton (2003) study the simplest version of the problem. Customers arrive with valuations drawn i.i.d. from an unknown distribution on $[0,1]$, and the seller posts a price from a continuous set. The demand curve $D(p)=\Pr(v\geq p)$ is unknown; the only feedback is whether each customer bought. The seller’s goal is to minimize regret against the single price $p^{*}$ that maximizes $p\cdot D(p)$. Kleinberg and Leighton prove that the minimax regret is $\Theta(\sqrt{T})$.187187187The upper bound, $O(\sqrt{T\log T})$, discretizes $[0,1]$ into $K=\lceil(T/\log T)^{1/4}\rceil$ prices and runs the UCB1 algorithm of Auer et al. (2002a). The lower bound, $\Omega(\sqrt{T})$, constructs a family of demand curves parameterized by the location of the optimal price $p^{*}\in[0.3,0.4]$. The key tension: posting prices far from $p^{*}$ is informative about demand but costly in revenue; posting prices near $p^{*}$ is cheap but uninformative. Resolving this tension costs at least $\Omega(\sqrt{T})$ in cumulative revenue. UCB1 (Auer et al., 2002a) selects the price maximizing $\hat{\mu}_{p_{k}}(t)+\sqrt{2\ln t/N_{p_{k}}(t)}$, where $\hat{\mu}_{p_{k}}(t)$ is the empirical mean profit and $N_{p_{k}}(t)$ the number of trials; the second term is an exploration bonus that shrinks as a price is tried more, implementing the principle of optimism in the face of uncertainty. In concrete terms, after 10,000 customers the seller loses roughly 100 customers’ worth of revenue to the uncertainty in demand. No algorithm can do better without imposing structure on $D$. For adversarial valuations (chosen by a worst-case opponent rather than drawn from a fixed distribution), the minimax regret rises to $\Theta(T^{2/3})$, achieved by the Exp3 algorithm188188188Exp3 (Auer et al., 2002b) maintains a weight $w_{k}(t)$ for each price, selecting $p_{k}$ with probability proportional to $w_{k}(t)$ and updating the chosen price’s weight by $\exp(\eta\hat{r}_{k,t})$ where $\hat{r}_{k,t}$ is the revenue importance-weighted by the selection probability; because no model of demand is assumed, the guarantee holds against an adversary who chooses valuations after observing the algorithm. on a discretized price grid. I focus on the stochastic setting throughout this chapter.